OpenStudy (anonymous):

what number must be added to both the numerator and the denominator of the fraction 5/21 to give a fraction equal to 3/7?

OpenStudy (king):

7

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

(5+x) / (21+x) = 3/7

OpenStudy (king):

let the number be x....then (5+x)/(21+x)=3/7 crossmultiply 35+7x=63+3x 4x=28 x=7

OpenStudy (anonymous):

let the number that you must add be called x. on the top you have 5+x while on the bottom you have 21+x together they must form a fraction equal to 3/7

OpenStudy (anonymous):

(5+x)/(21+x)=3/7 crossmultiply 35+7x=63+3x 4x=28 x=7

OpenStudy (anonymous):

^Quit cheating -_-

OpenStudy (anonymous):

fine then

OpenStudy (anonymous):

Just stop. -.-

OpenStudy (anonymous):

why

OpenStudy (anonymous):

u stop

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

now all we need is a "let me wolframalpha that for you" site